Lift After Gamma Mods for Cintel MK3
- 1 LIFT AFTER GAMMA
LIFT AFTER GAMMA
As created and written by Dave Corbitt
Modification and alignment for Analog Color Channel of Rank Cintel MK3 series telecines.
INSTALL AT OWN RISK
Please note the revision to the modifications in the 100251 board below
The MK3 series have a built in limiting factor for video level adjustments during Negative mode colorgrading. This problem is due to how and where the various reference clamp pulses are inserted in the processing chain of the Colorgrade. Adjusting LIFT causes a corresponding change in WHITE levels in the same direction. The telecine operator must then readjust GAIN to compensate. This is a problem for at least three reasons;
1. "Rides" on "One Lights" become difficult.
2. Correction Range is limited. This then becomes a problem on dailies with wide exposure ranges.
3. Color Correction sessions take too long due to control interactions.
With the enclosed modification incorporated in to the Color Channel, these problems are resolved. Note; this mod is not appropriate for Enhanced MK3C's that have the PEC controls interconnected to the Colorgrade controls. All other MK3 series machines however will benefit including various versions of the Turbo . Later versions of the Turbo have already incorporated an earlier version of this modification and will perform similarly to a modified machine as described in this update. Note: All new resistors should be 2% or better and low noise (metal film) types.
MODS TO VIDEO RACK MOTHERBOARD
1. Add three wires to the motherboard
2. Connect Red 100247 pin 33 to Red 100249 pin 2.
3. Do the same for Green and Blue.
MODS TO ALL THREE 100247'S LOG AMPS (RED, GREEN, BLUE)
1. Remove VT2. Connect PCB pads for VT2 C to E with small wire links.
2. Remove R11 and MR5. Replace MR5 with wire link. Leave R11 open.
3. R73. Change from 1.5k to 4.7k low noise high precision type (1 or 2%, metal film).
MODS TO ALL THREE 100249'S EXP AMPS (RED, GREEN, BLUE)
1. Remove R55 link. Install 6k8.
2. R5. Install 2k7.
3. R60. Install 15k.
4. C10. Install 0.1 uF/ 50 v.
MODS TO 100251
Note revision below
1. Remove VT 7, 12, and 17. Use 3 small wire links to connect E to C on pc pads of those transistor positions.
2. Remove R 1, 2, 3, and 9.
3. Isolate the collectors of VT 9, 14, and 19 by cutting the traces. Restore any connections broken while doing this other than to the above transistors. Reconnect the Collectors of VT 9, 14, and 19 to IC 6 pin 6.
4. Reconfigure IC 1 as in Fig. 1 below. Also, disconnect MON TP4 (on front cover plate of the board) from R94/93 junction and reconnect to IC 1 pin 6 through a 1k resistor as in sketch below ( Fig 1). Some traces must be cut to do this. Compare this sketch to original drawing for 100251.
The color channel should remain aligned as before except for the following;
1. Measure the voltage difference between 100251 MON TP4 ( on 100251 front panel) and 100252 REF TP7 ( on 100252 front panel). Adjust RV5 on 100251 for minimum difference.
2. With Analog Test Ramp ON check black clip levels are correct on 100249 outputs. If incorrect, adjust RV3 on each 249. Select a new value for R60 if RV3 has insufficient range. Master lift should be at reference setting while setting black clip.
3. With POS mode and TEST both ON, adjust Master Lift so Test Tamp is just above black clip. Engage Negative Mode. the inverted Test ramp should now have its negative peak just above black clip also. If not, readjust RV5 on 100251 to achieve this. (See FIG. 2 below).
edited by David Tosh